Find the level curve each of function at indicated point? (a) 𝒇(𝒙, π’š) = 𝒙 𝟐 βˆ’ π’š 𝟐 , 𝒂𝒕 𝒄 = 𝟏. (b) π’ˆ(𝒙, π’š) = 𝒆 π’™π’š , 𝒂𝒕 𝒄 = 𝟎. πŸ“, πŸ‘. (c) 𝒉(𝒙, π’š) = 𝒍𝒏 (𝒙 𝟐 + π’š 𝟐 ) 𝒂𝒕 𝒄 = βˆ’πŸ, 𝟎, 𝟏.Β 


1
Expert's answer
2022-01-18T06:32:34-0500

"a)x^2-y^2=1\\
b)e^{xy}=0\\
C)In(x^2+y^2)=-1\\
In(x^2+y^2)=0\\
C)In(x^2+y^2)=1"